Catch-up photos:
- Continuing to work on removing TBA bolts.
- The bearing unit was finally removed, and stowed in a bag.
- After that, the mounting unit was un-bolted from the inboard race ring and also stowed in a bag.
- The race ring with the TBA fully removed.
The next task in the TBA R&R is to lubricate the ring, and that's a one-person job, so Moghbeli is going to head up to the Camera Port 8 camera to reposition a cable that's obstructing some motion of the camera assembly.
Grease gun and dry wipe. Also, sun's out again!
Moghbeli's view at the CP8 worksite. Ground crews are discussing how best to secure that cable.
Three clicks of grease applied with the J-hook - McClain approvingly notes that O'Hara's gloves are still white.
Looking over Moghbeli's shoulder from a station cam as she works on the CP8 Ethernet cable.
Has someone the EMU numbers (not the SAFER numbers)?
Greasing the other two surfaces of the race ring (flat and outer canted) with the straight grease gun.
Moghbeli is going to demate the cable, put a twist in it, and re-mate. (Once it gets unstuck, that is.)
Edit - fun fact, that bar thing with the two grapple fixtures on the radiator unit was the first un-pressurized cargo to be delivered to station in a Dragon trunk, back on CRS-2.
EV1 is still working on that camera group. EV2 has started one bolt on the mounting bracket by hand - no HECA views of that, though, I think they mentioned her camera was running a bit hot and were going to turn it off for a sec.
Edit - there's a camera view! You can also see a good amount of grease on the race ring when the sun's at the right angle.
Good views of the bearing package, which O'Hara has slid into the mounting bracket.
O'Hara is working to drive Bolt 2 on the TBA, which will cause the back portion to move down and clamp around the triangular race ring. Seems to be drifting around a lot, though.
Ground controllers have decided that they will not bring the RFG in the airlock on today's EVA, but they will remove the wedge clamps from the back edge.
O'Hara put about 11 turns into Bolt 2 - less than expected, but it's torquing out when she tries to do another turn. Mission Control seems OK with it.
Bolts 1 and 3 have also been driven.
More bolt torquing ops - didn't get quite as many turns on bolt 3 as expected.
Caps off the cable and socket, inspection looks good.
Cable has been connected.
Moghbeli is trying to track down a crew lock bag that she dropped off earlier in the spacewalk.